Why generally we take exponent of 10 same as that of 5 but ninot 2?how to decide exponent of any number in factorial notation.with proof

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

It can be understood in terms of limited reagent concept

for ex
10 = 5^1 * 2^1
20 = 5^1 * 2^2
similarly 1000 = 5^3 * 2^3
but 2000 = 5^3 * 2^4

so whatever is in lower number we can treat it as a controlling factor for representing a number as 10 power

Regards